### v1

@@ 2.2 Modeling choices for the core kernel
In principle, there are no constraints regarding statistical modeling of the core kernel; any positive semi-definite function can be used. In contrast to many approximate methods to scale GPs, our framework allows us to use the full suite of tools that have been developed for kernel modeling, including stationary and nonstationary kernels. For example, even using standard stationary kernels (e.g., the Matérn class) as the core kernel results in the overall kernel being both nonstationary and sparse (since $C_{\text{sparse}}$ is itself nonstationary). Later in the paper, we utilize parametric families of nonstationary kernels [drawn from, e.g., 15, 44, 45]; in principle, however, one could use any other nonstationary kernel including deep kernels [17] or deep GPs [16]. In the absence of specific knowledge regarding the form of the core kernel, “parametric” nonstationary kernels [46] such as those derived in [15] and later extended in [45] have been shown to perform well in practice. See Appendix B for more details.
2.3 Special cases
@@
Prior specification and regularization allow the kernel to learn the underlying true state (see Appendix D.2).
@@ 3 Bayesian stochastic modeling and inference
We now propose a Bayesian framework for statistical modeling with a Gaussian process (Section 3.1) along with an approach to posterior prediction with uncertainty quantification (Section 3.2). Details regarding the computational requirements of our method are discussed in Section 3.3, and a pair of synthetic data tests conducted to verify the performance of our methodology (described in Appendix E) are summarized in Section 3.4.
@@ 3.1 Bayesian GP with prior specification, regularization, and training
where $\varepsilon(\cdot)$ is a stochastic error component that is independently distributed as ${N}(0,\tau^{2}({\bf x}))$, with $\varepsilon(\cdot)$ and $y(\cdot)$ are independent. We furthermore assume that the error variance process $\tau^{2}(\cdot)$ is known up to hyperparameters ${\boldsymbol{\theta}}_{z}$. The stochastic process $y(\cdot)$ is modeled as a Gaussian process, where the prior mean function $E[y({\bf x})]={\bf w}({\bf x})^{\top}{\boldsymbol{\beta}}$ is linear in a set of covariates or basis functions and the kernel $C_{y}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{y})\equiv Cov\big{(}y({\bf x}),y({\bf x}^{\prime})\big{)}$ is assumed known up to hyperparameters ${\boldsymbol{\theta}}_{y}$ and describes the covariance between the process $y(\cdot)$ for all ${\bf x},{\bf x}^{\prime}\in\mathcal{X}$.
@@
[…]
@@
$C_{z}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}})=C_{y}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{y})+\tau({\bf x})\tau({\bf x}^{\prime})I_{\{{\bf x}={\bf x}^{\prime}\}},\hskip 17.22217pt\text{for all }{\bf x},{\bf x}^{\prime}\in\mathcal{X},$ (4)
@@
$s_{0}\sim U(0,10^{5});\hskip 21.52771ptr_{0}\sim U(0,D_{0});\hskip 21.52771ptr_{ij}\sim U(0,D_{r}),\hskip 8.61108pti=1,\dots,n_{1};j=1,\dots,n_{2}$
@@
For GPs with trained hyperparameters, the posterior distribution (5) is not available in closed form regardless of prior choice, and so we must resort to Markov chain Monte Carlo (MCMC) methods to conduct inference on ${\boldsymbol{\beta}}$ and ${\boldsymbol{\theta}}$ [47]. See Appendix C for further details.
@@ 3.2 Posterior prediction with uncertainty quantification
[…]
@@ 3.3 Computational considerations
In order to utilize $C_{z}$ within a Bayesian framework, each iteration of the Markov chain Monte Carlo for the posterior Eq. 5 requires at least one evaluation of the log of the probability density function given in Eq. 3 (the log-likelihood), i.e., $\log p({\bf z}|{\boldsymbol{\beta}},{\boldsymbol{\theta}})\propto-0.5\log\det({\bf C}_{z})-0.5({\bf z}-{\bf W}{\boldsymbol{\beta}})^{\top}{\bf C}_{z}^{-1}({\bf z}-{\bf W}{\boldsymbol{\beta}})$, where the elements of ${\bf C}_{z}$ come from Eq. 4. From a computational standpoint, there are three main tasks associated with this evaluation: (1) calculating the covariance matrix ${\bf C}_{z}$, (2) calculating the log-determinant $\log\det({\bf C}_{z})$, and (3) solving the quadratic form $({\bf z}-{\bf W}{\boldsymbol{\beta}})^{\top}{\bf C}_{z}^{-1}({\bf z}-{\bf W}{\boldsymbol{\beta}})$. The sparsity-discovering nature of our kernel means there are no shortcuts for calculating the covariance matrix for a fixed set of hyperparameters: different elements of ${\bf C}_{z}$ will be zero vs. nonzero for different hyperparameter configurations. As such, we must first calculate ${\bf C}_{z}$ in dense form in batches that are large enough to utilize the chosen hardware well. Computationally, the covariance batches are dense but contain many zeros because of our kernel designs; therefore, each batch is cast to a sparse format on the worker before being transferred back to the host process. There, the full covariance matrix is assembled and exists only in a sparse format leading to minimal RAM requirements. For particularly large datasets, the computation of the batches can be performed by utilizing distributed high-performance-computing architecture. For example, [40] applied a related kernel to a climate data set of more than 5 million points using HPC resources at the National Energy Research Scientific Computing Center at Lawrence Berkeley National Laboratory.
@@
Steps two and three, namely the log-determinant calculation and the linear-system solve related to the quadratic form can be performed at low computational costs due to the sparse covariance matrix. For example, using the Lanczos method to estimate the log-determinant has a time complexity of $\mathcal{O}(km+k^{2})$, where $m$ is the number of nonzero entries in the matrix and $k$ is the number of iterations, which depends on the eigenvalue distribution of the matrix. Further, the conjugate gradients methods to solve linear systems has time complexity $\mathcal{O}(mk)$, where $m$ is the number of nonzero elements in the matrix and $k$ is the number of iterations which depend on the condition number of the matrix. For poorly conditioned covariance matrices, we use the minimum residual method which has the same time complexity but is less sensitive to the condition number. In practice, we frequently see sparsity numbers of $1/10000$ for large problems of 1 million data points which, together with the proposed kernel design, leads to log-determinant and linear solve compute times in the order of a few second. For smaller problems ($<100,000$), we can take advantage of LU or Cholesky decompositions.
@@
The fact that our methodology yields no shortcuts for calculating ${\bf C}_{z}$ is in sharp contrast with approximate methods, e.g., the nearest neighbor Gaussian process (NNGP) proposed in [48] and the Vecchia approximations proposed in [32]. First, the NNGP and Vecchia approximations imply a sparse precision matrix (i.e., inverse covariance), while our method implies a sparse covariance. Second, the implicit conditioning structure of each method implies that for a given ordering of points and method for choosing nearest neighbors [see choices C1-C5 given in Section 2.3 of 32], some elements of the precision matrix will always be nonzero, and other elements will always be exactly zero. The fixed nonzero elements are known ahead of time (again, based on specific choices made for ordering of points and selecting neighbors), such that likelihood evaluations for NNGP and Vecchia never involve dense matrices. At the end of the day, our methodology is still an exact Gaussian process: some of the usual challenges associated with computation for a GP cannot be avoided. However, what we gain is substantially less subjectivity: no a priori choices need to be made regarding, e.g., the number of neighbors to condition on, the ordering of points, and the selection of nearest neighbors.
3.4 Synthetic data tests
Before applying our methodology to a real-world data set, we conducted two synthetic data tests to ensure that our proposed kernel is robust to mis-specification of the number of bump functions and performs well relative to competing methods across a variety of ground-truth data states.
Our first test (see Appendix D for complete details) verifies that our proposed sparse kernel can correctly identify the bump function placement and size, and furthermore, the regularization method for the bump functions described in Section 3.1 effectively eliminates the effect of any bump functions that may be unnecessary. For a real data example, this leads us to recommend specifying more bump functions than one might think is truly necessary (i.e., err on the side of making $n_{1}$ and $n_{2}$ large). Of course, more bump functions means more hyperparameters to learn during training, so in practice, one must maintain a balance between too many hyperparameters and choosing $n_{1}$ and $n_{2}$ to be “large enough.”
Our second test verifies that our kernel performs as well or better than competing kernels within a GP across a variety of ground truth data-generating mechanisms in terms of root mean square error (RMSE) and the continuous rank probability score [CRPS; 49] for out-of-sample test data. This assessment involves first generating data from a mean-zero Gaussian process with a specified kernel. Four kernels are considered: non-sparse and stationary (denoted “scenario” 1 or S1), sparse and stationary (denoted S2), sparse and nonstationary (denoted S3), and non-sparse and nonstationary (denoted S4). See Appendix E for more details and sample draws from each GP. For S1-S4, the synthetic data drawn from the parent distribution represent a ground truth function; we generate $100$ draws, each of which has $50$ (randomly sampled) training measurements and $300$ (regularly spaced) test data points for validation. Each training data set includes the addition of white noise, which is approximately 10% of the signal variance. For each scenario and each synthetic draw, we fit a series of constant-mean GP models to the training data, mirroring the example given in Figure 1: a stationary Matérn kernel (“model” 1 or M1), the sparse general Vecchia (SGV) approximate GP method [32] with a stationary Matérn kernel (M2), our sparsity-discovering kernel proposed from Section 2 with a stationary Matérn core kernel (setting $n_{1}=4$ and $n_{2}=4$; M3), and the same sparsity-discovering kernel with the nonstationary core kernel proposed in Equation B.1 (M4). See Appendix E for further details on training the various Gaussian process models.

### v3

@@ 2.2 Modeling choices for the core kernel
In principle, there are no constraints regarding statistical modeling of the core kernel; any positive semi-definite function can be used. In contrast to many approximate methods to scale GPs, our framework allows us to use the full suite of tools that have been developed for kernel modeling, including stationary and nonstationary kernels. For example, even using standard stationary kernels (e.g., the Matérn class) as the core kernel results in the overall kernel being both nonstationary and sparse (since $C_{\text{sparse}}$ is itself nonstationary). Later in the paper, we utilize parametric families of nonstationary kernels [drawn from, e.g., 17, 45, 46]; in principle, however, one could use any other nonstationary kernel including deep kernels [19] or deep GPs [18]. In the absence of specific knowledge regarding the form of the core kernel, “parametric” nonstationary kernels [47] such as those derived in [17] and later extended in [46] have been shown to perform well in practice. See Appendix B for more details.
2.3 Special cases and infill asymptotics
@@
Prior specification and regularization allow the kernel to learn the underlying true state (see Appendix 4.2).
Regardless of which case emerges from a given application, we can derive the asymptotic properties our proposed method under the infill framework. As the in-fill design becomes dense, the posterior concentrates on the true kernel parameters, including the bump function radii that dictate where the covariance vanishes. Under this regime, we can establish the consistency of a GP using classical results like Theorem 1 of [48], those radii lock onto fixed finite values, so every row of the covariance matrix carries only a bounded number of non-zero entries regardless of sample size. Once this data-driven support has stabilized, the model behaves exactly like a Gaussian process with a conventional, (additive) compactly supported kernels, allowing all classical consistency and efficiency results for such kernels to apply unchanged. We defer a formal statement of this result and the proof to Appendix A.
@@ 3 Bayesian stochastic modeling and inference
We now propose a Bayesian framework for statistical modeling with a Gaussian process (Section 3.1) along with an approach to posterior prediction with uncertainty quantification (Section 3.2). Details regarding the computational requirements of our method are discussed in Section 3.3.
@@ 3.1 Bayesian GP with prior specification, regularization, and training
where $\varepsilon(\cdot)$ is a stochastic error component that is independently distributed as ${N}(0,\tau^{2}({\bf x}))$, with $\varepsilon(\cdot)$ and $y(\cdot)$ are independent. We furthermore assume that the error variance process $\tau^{2}(\cdot)$ is known up to hyperparameters ${\boldsymbol{\theta}}_{z}$. The stochastic process $y(\cdot)$ is modeled as a Gaussian process, where the prior mean function $E[y({\bf x})]={\bf w}({\bf x})^{\top}{\boldsymbol{\beta}}$ is linear in a set of covariates or basis functions and the kernel $C_{y}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{y})\equiv Cov\big(y({\bf x}),y({\bf x}^{\prime})\big)$ is assumed known up to hyperparameters ${\boldsymbol{\theta}}_{y}$ and describes the covariance between the process $y(\cdot)$ for all ${\bf x},{\bf x}^{\prime}\in\mathcal{X}$. Note that for spatially-indexed data, $\mathcal{X}=\mathcal{S}\subset\mathbb{R}^{2}$ (assuming two dimensional coordinates, e.g., longitude/latitude); for spatio-temporal data where the time dimension is assumed to be continuous, $\mathcal{X}=\mathcal{S}\times\mathcal{T}\subset\mathbb{R}^{3}$.
@@
$p({\bf z}|{\boldsymbol{\beta}},{\boldsymbol{\theta}})=\int p({\bf z}|{\bf y},{\boldsymbol{\theta}}_{z})p({\bf y}|{\boldsymbol{\beta}},{\boldsymbol{\theta}}_{y})d{\bf y}={N}\big({\bf W}{\boldsymbol{\beta}},{\bf\Delta}({\boldsymbol{\theta}}_{z})+{\bf\Omega}({\boldsymbol{\theta}}_{y})\big),$ (3)
@@
$C_{z}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}})=C_{y}({\bf x},{\bf x}^{\prime};{\boldsymbol{\theta}}_{y})+\tau({\bf x})\tau({\bf x}^{\prime})I_{\{{\bf x}={\bf x}^{\prime}\}},\qquad\text{for all }{\bf x},{\bf x}^{\prime}\in\mathcal{X},$ (4)
@@
$s_{0}\sim U(0,10^{5});\hskip 21.52771ptr_{0}\sim U(0,D_{0});\hskip 21.52771ptr_{ij}\sim U(0,D_{r}),\quad i=1,\dots,n_{1};j=1,\dots,n_{2}$
@@
Despite the fact that the likelihood and $p({\boldsymbol{\beta}})$ are Gaussian, the fact that the kernel parameters $\boldsymbol{\theta}$} are unknown means that the posterior distribution (5) is not available in closed form regardless of prior choice $p({\boldsymbol{\theta}}$). Therefore, we must resort to Markov chain Monte Carlo (MCMC) methods to conduct inference on $\boldsymbol{\beta}$} and $\boldsymbol{\theta}$} [49]. See Appendix C for further details.
@@ 3.2 Posterior prediction with uncertainty quantification
$p({\bf y}_{P}|{\bf z})=\int p({\bf y}_{P},{\boldsymbol{\beta}},{\boldsymbol{\theta}}|{\bf z})d{\boldsymbol{\beta}}d{\boldsymbol{\theta}}=\int p({\bf y}_{P}|{\boldsymbol{\beta}},{\boldsymbol{\theta}},{\bf z})p({\boldsymbol{\beta}},{\boldsymbol{\theta}}|{\bf z})d{\boldsymbol{\beta}}d{\boldsymbol{\theta}}.$
@@ 3.3 Computational considerations
In order to utilize $C_{z}$ within a Bayesian framework, each iteration of the Markov chain Monte Carlo for the posterior Eq. 5 requires at least one evaluation of the log of the probability density function given in Eq. 3 (the log-likelihood), i.e., $\log p({\bf z}|{\boldsymbol{\beta}},{\boldsymbol{\theta}})\propto-0.5\log\det({\bf C}_{z})-0.5({\bf z}-{\bf W}{\boldsymbol{\beta}})^{\top}{\bf C}_{z}^{-1}({\bf z}-{\bf W}{\boldsymbol{\beta}})$, where the elements of ${\bf C}_{z}$ come from Eq. 4. From a computational standpoint, there are three main tasks associated with this evaluation: (1) calculating the covariance matrix ${\bf C}_{z}$, (2) calculating the log-determinant $\log\det({\bf C}_{z})$, and (3) solving the quadratic form $({\bf z}-{\bf W}{\boldsymbol{\beta}})^{\top}{\bf C}_{z}^{-1}({\bf z}-{\bf W}{\boldsymbol{\beta}})$. The sparsity-discovering nature of our kernel means there are no shortcuts for calculating the covariance matrix for a fixed set of hyperparameters: different elements of ${\bf C}_{z}$ will be zero vs. nonzero for different hyperparameter configurations. In other words, we cannot know ahead of time which elements are zero, and hence must consider all pairs of points when calculating ${\bf C}_{z}$. To get around this bottleneck, instead of calculating a very large covariance matrix on a single node, we first calculate ${\bf C}_{z}$ in dense form in batches that are large enough to utilize the chosen hardware well: the “batches” represent blocks of the full covariance matrix that can be calculated in parallel (note that the use of batches does not introduce any approximations into the calculation). Computationally, the covariance batches are dense but contain many zeros because of our kernel design; therefore, each batch is cast to a sparse format (i.e., we store only the non-zero elements and their row-column position) on the worker before being transferred back to the host process. There, the full covariance matrix is assembled and exists only in a sparse format leading to minimal RAM requirements. For particularly large datasets, the computation of the batches can be performed by utilizing distributed high-performance-computing architecture. For example, [50] applied a related kernel to a climate data set of more than 5 million points using HPC resources at the National Energy Research Scientific Computing Center at Lawrence Berkeley National Laboratory.
@@
Steps two and three, namely the log-determinant calculation and the linear-system solve related to the quadratic form can be performed at low computational costs due to the sparse covariance matrix. For example, using the Lanczos method to estimate the log-determinant has a time complexity of $\mathcal{O}(km+k^{2})$, where $m$ is the number of nonzero entries in the matrix and $k$ is the number of iterations, which depends on the eigenvalue distribution of the matrix. In the biggest data examples in Section 5.2, we used the Imate Python package’s Lanczos method with default settings: a normally distributed error in the log-determinant of about 0.1% (the impact of this on the MCMC’s acceptance probability was negligble). The conjugate gradients (CG) method to solve linear systems has time complexity $\mathcal{O}(mk)$, where $m$ is the number of nonzero elements in the matrix and $k$ is the number of iterations which depend on the condition number of the matrix. Unfortunately, in practice, since our proposed kernel results in a possibly large range of eigenvalues for the covariance matrix, we found that CG frequently failed to converge. Instead, we used the minimum residual (MINRES) method which has the same time complexity but is more robust to wildly varying eigenvalues and converges well. We always initialized the MINRES algorithm randomly and with a set error tolerance of $10^{-8}$. In practice, we frequently see sparsity numbers of $1/10000$ for large problems of 1 million data points which, together with the proposed kernel design, leads to log-determinant and linear solve compute times in the order of a few seconds. Ultimately, in our experience, it has turned out that the combination of the MINRES for linear solves and the Lanczos method for the log-determinant was the only viable option for applying our kernel to data sets with more than one million points. For smaller problems (<100,000<100,000), we take advantage of LU or Cholesky decompositions.
@@
The fact that our methodology yields no shortcuts for calculating ${\bf C}_{z}$ is in sharp contrast with approximate methods, e.g., the nearest neighbor Gaussian process (NNGP) proposed in [51] and the Vecchia approximations proposed in [33]. First, the NNGP and Vecchia approximations imply a sparse precision matrix (i.e., inverse covariance), while our method implies a sparse covariance. Second, the implicit conditioning structure of each method implies that for a given ordering of points and method for choosing nearest neighbors [see choices C1-C5 given in Section 2.3 of 33], some elements of the precision matrix will always be nonzero, and other elements will always be exactly zero. The fixed nonzero elements are known ahead of time (again, based on specific choices made for ordering of points and selecting neighbors), such that likelihood evaluations for NNGP and Vecchia never involve dense matrices.
To explore a more detailed runtime analysis of the computational cost associated with our method, we conducted a test using synthetic data across different sample sizes ($N\in\{50000,100000,200000\}$), choices of $n_{1}$ and $n_{2}$ ($n1\times n2\in\{2\times 50,4\times 25,10\times 10\}$), and sparsity level (ranging from 1e-05 to 1e-01). For each of these combinations, we separately tallied the time needed to calculate the covariance ${\bf C}_{z}$, the log-determinant calculation using the Lanczos method, and the MINRES linear-system solve; see Figure 3. The covariance compute times are based on using 16 GPUs to parallelize across covariance batches; the Lanczos and MINRES times correspond to a single GPU. Generally speaking, the total time needed to evaluate the likelihood ranges from approximately 2 seconds for $N=50000$ up to about 10 seconds for $N=200000$. The slowest operations are calculation of the covariance and the log-determinant, while the MINRES solve is very fast regardless of sample size, $n_{1}\times n_{2}$ configuration, and sparsity. As we would expect, the computational time for calculating the covariance is unaffected by the sparsity: by definition, the covariance calculation requires querying all pairs of points to determine which entries are zero versus non-zero. The computational time is also largely unaffected by the $n_{1}\times n_{2}$ configuration: for a given sample size/sparsity combination, the compute time is roughly the same across $2\times 50$, $4\times 25$, and $10\times 10$. Interestingly, the MINRES and log-determinant calculation are somewhat insensitive to the sparsity, at least in cases where the fraction of nonzero entries is less than 1e-03.
Figure 3: Runtime analysis of the computational cost associated with the proposed methodology. We show the compute time needed for three calculations: (1) the covariance matrix ${\bf C}_{z}$; (2) the log-determinant using the Lanczos method; and (3) the MINRES linear system solve. Compute time is shown across different combinations of the sample size (rows), the $n_{1}\times n_{2}$ configuration (columns), and sparsity level ($x$-axis).
In summary, we reiterate that our methodology is still an exact Gaussian process: the implication is some the computational bottlenecks associated with exact GPs cannot be avoided. We make no claims about the computational efficiency of our method relative to approximate GP methods like Vecchia for moderately sized data sets. However, the benefits are clear: we avoid any approximations or alternative formulations (e.g., assuming sparse precision matrices or autoregressive structure) while developing a highly flexible kernel that, when combined with HPC resources, can be used for fully Bayesian analysis of large data sets, on the order of millions of data points. Furthermore, we also gain substantially less subjectivity: no a priori choices need to be made regarding, e.g., the number of neighbors to condition on, the ordering of points, and the selection of nearest neighbors.
4 Synthetic data examples
We now describe two synthetic data examples to demonstrate the performance of our proposed kernel. The first evaulates performance under mis-specification of the number of bump functions and the efficacy of our proposed regularization priors when too many bump functions are permitted (Section 4.1). The second compares our proposed methodology with competing methods across a variety of ground-truth data states (Section 4.2). In Section 4.2, we assess performance quality using two metrics: root mean square error (RMSE) for the test set, which assesses the pointwise error in posterior mean predictions; and the continuous rank probability score [CRPS; 52] for the test set, which (unlike the RMSE) assesses both precision and sharpness of predictions. For example, the CRPS rewards a bad prediction with large uncertainty over a bad prediction with small uncertainty. Note that the reported CRPS scores are calculated for each test data point and then averaged across all test data.
4.1 Mis-specified bump functions
and we have noisy measurements from this function for training a GP (see Figure 4a.). The form of this ground truth function suggests that there should be two sets of two bump functions: $y(x)$ is perfectly correlated in $(0,0.25]$ and $(0.5,0.75]$; furthermore, $y(x)$ is perfectly correlated in $(0.25,0.5]$ and $(0.75,1)$. In other words, $n_{1}=2$ and $n_{2}=2$. We then fit our proposed kernel to the noisy data assuming we know nothing about its ground truth, under three configurations of bump functions: too few bump functions, where we set $n_{1}=1$ and $n_{2}=2$; the right number of bump functions, using $n_{1}=2$ and $n_{2}=2$; too many bump functions, using $n_{1}=4$ and $n_{2}=4$. Notice that although the second configuration is the only one with the correct number of bump functions, in all cases we infer the bump functions’ characteristics (positions and radii) directly from the data. When fitting, we suppose the core kernel is a stationary Matérn with smoothness $\nu=2.5$.
Figure 4: Results for the synthetic data test described in Sections 4.1. The ground truth function and noisy measurements are shown in panel (a); panels (b) and (c) show the best estimates of the bump functions (posterior medians) and posterior predictions with uncertainty (posterior mean and 90% credible intervals), respectively. Results are shown when the kernel has too few bump functions (left), the correct number of bump functions (center), and too many bump functions (right).
Results are shown in Figure 4(b) and (c), where we show both the posterior median estimates of the bump functions and the posterior predictions of $y(\cdot)$ with uncertainty, respectively. First, consider the bump function estimates in Figure 4(b). When there are too few bump functions, the training correctly identifies the position and size of the bump functions for part of the domain, i.e., that $(0.25,0.5]$ is correlated with $(0.75,1)$, while the kernel does not have the flexibility to capture the distance-unrelated correlations in $(0,0.25]$ and $(0.5,0.75]$. As such, the predictions capture the ground truth function with very low uncertainty in $(0.25,0.5]$ and $(0.75,1)$ but revert to the sparse stationary kernel in the other parts of the domain – effectively over-fitting the data in $(0,0.25]$ and $(0.5,0.75]$. When there are the correct number of bump functions, the training correctly identifies where these bump functions should go and how large they should be; consequently, posterior predictions are correctly calibrated over the entire unit interval. Finally, when there are too many bump functions, the algorithm correctly determines both the size and position of the bump functions, but also that only four of the 16 total bump functions should be activated. The other 12 are appropriately zeroed out via regularization to ensure that we avoid both under- and over-fitting. Posterior predictions are effectively identical to those generated using the correct number of bump functions.
4.2 Comparison with competing methods
  1. S1: Non-sparse and stationary: $C_{y}$ is the isotropic Matérn kernel.
  2. S2: Sparse and stationary: $C_{y}({\bf x},{\bf x}^{\prime})=\sigma^{2}f_{0}({\bf x},{\bf x}^{\prime};r_{0})$, where $f_{0}$ is as in Eq. 2.1.
  3. S3: Sparse and nonstationary: $C_{y}({\bf x},{\bf x}^{\prime})=\sigma({\bf x})\sigma({\bf x}^{\prime})f_{0}({\bf x},{\bf x}^{\prime};r_{0})$, with input-dependent $\sigma(\cdot)$.
  4. S4: Non-sparse and nonstationary: here we take $C_{y}$ to be the nonstationary kernel described in [46].
Here we use S1-S4 to denote different “scenarios”; see Appendix D for more details on the data generating mechanisms, specific hyperparameters used, and sample draws from the true parent distribution. In each case, the synthetic data drawn from the parent distribution represents a ground truth function; we can therefore generate $N_{\text{train}}=50$ (randomly sampled) and $N_{\text{test}}=300$ (regularly spaced) data points for validation. We also explore the effect of how large the noise is relative to the signal, and add white noise to the ground truth with a variance that is 5% (signal-to-noise ratio or SNR of 20), 10% (SNR = 10), and 20% (SNR = 5) of the signal variance. For each scenario, we generate $N_{\text{rep}}=50$ replicates or draws from the ground truth distribution; note that the locations of the training data are different across each replicate.
For each scenario and each replicate, we fit a series of constant-mean GP models to the synthetic data, summarized in Table 1 and mirroring the example given in Figure 1. As a baseline, we fit a stationary Gaussian process with a Matérn kernel, denoted Model 1 or “M1”. To compare against this baseline, we consider a suite of GP methods: the Vecchia approximation [M2, 33]; the predictive process [M3, 26]; covariance tapering [M4, 30]; our proposed approach, using the kernel proposed in Section 2 with a stationary Matérn core kernel (M5); and as in M5, but with a nonstationary core kernel defined in Equation B.1 (M6). The Vecchia approximation uses 15 nearest-neighbors; the predictive process uses 12 knots ($12/50=24\%$ of the data size); and the prior for the covariance taper limits the nonzero correlations to be within $\pm 1$ unit ($10\%$ of the size of the input space) of each data point. In all cases, the proposed covariance kernel is used within a fully Bayesian model where all hyperparameters are trained separately for each replicate.
In principle, M1-M4 could involve nonstationary kernels; however, we maintain stationary kernels for two reasons. First, M1 is primarily used here as a reference for what one can obtain with a “default” (stationary and exact) GP when working with smaller data sets; the whole point of this methodology is that M1 will become unusable for much larger data sets. Second, all of the underlying theory for Vecchia approximations, predictive process, and covariance tapering has been derived for stationary kernels. For example, one can use the Vecchia framework with a nonstationary kernel [see, e.g., 46]; however, without further theoretical developments there are no guarantees that the subjective choices one must make for the Vecchia approximation (size of conditioning sets, ordering of points, selection of nearest neighbors) will result in an appropriate approximation for a nonstationary kernel. Our implementation of M1 and M2 involve standard, off-the-shelf implementations of these methods as enabled by the BayesNSGP software package for R [53]. We implement our own versions of M3 and M4, since existing software [e.g., 54] cannot be used for one-dimensional inputs. For M5 and M6, the main user-specified feature is determining how many bump functions are allowed: here, we set $n_{1}=4$ and $n_{2}=4$). In order to ease comparison, we show CRPS for M2-M6 relative to corresponding quantities from M1: relative scores less than one indicate CRPS is smaller (and therefore better) for the model of interest versus a stationary GP.
Data generating mechanisms Kernels within the Gaussian process
Label Description Label Description
S1 Non-sparse, stationary M1 Stationary Matérn kernel
S2 Sparse, stationary M2 Vecchia approximation with Matérn
S3 Sparse, nonstationary M3 Predictive process
S4 Non-sparse, nonstationary M4 Covariance tapering
M5 Sparsity-discovering kernel with stationary core
M6 Sparsity-discovering kernel with nonstationary core
Table 1: Summary of five data-generating mechanisms (or “scenarios”) used to generate synthetic data and six kernels used within a Gaussian process to fit each of $N_{\text{rep}}=50$ repeated draws from each scenario.